OpenStudy (iimanii):

What is the proportion? 6/34 = p/68

OpenStudy (study_buddy99):

what times 34 is 68?

OpenStudy (iimanii):

22,312

OpenStudy (iimanii):

2,312

OpenStudy (study_buddy99):

good 2 times 34 is 68

OpenStudy (iimanii):

yes

OpenStudy (study_buddy99):

so 6 times 2 is what?

OpenStudy (iimanii):

12

pooja195 (pooja195):

6/34 = p/68 simply cross multiply \[\huge~\rm~\bf~\frac{ 6 }{ 34 }=\frac{ p }{ 68 }\] 34*p=? 68*6=? @study_buddy99 the way you are doing it is incorrect where did the 2 come from ?

OpenStudy (iimanii):

im confused

OpenStudy (study_buddy99):

34*2=68 so 6*2=12

OpenStudy (iimanii):

SO ITS 12?

OpenStudy (study_buddy99):

yes

OpenStudy (iimanii):

opps i put my caps on lol

OpenStudy (iimanii):

ok

pooja195 (pooja195):

You want to figure out the proportion using cross multiplcation @study_buddy99 did the work for you by giving you the answer. And you probably don't understand. .

OpenStudy (iimanii):

can you guys help me with other question

OpenStudy (iimanii):

ok can you explain to me how it goes then

pooja195 (pooja195):

Do you understand how @study_buddy99 got 2?

OpenStudy (iimanii):

yes

pooja195 (pooja195):

How did they get 2?

OpenStudy (iimanii):

because 32 times what equals 68 (2)

OpenStudy (iimanii):

32 times2 equals 68

pooja195 (pooja195):

You know what's funny? 32 isn't even part of your problem.

OpenStudy (iimanii):

i ment 34 lol

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @study_buddy99 what times 34 is 68? \(\color{#0cbb34}{\text{End of Quote}}\) Question asked \(\color{#0cbb34}{\text{Originally Posted by}}\) @IImanii 22,312 \(\color{#0cbb34}{\text{End of Quote}}\) Your response.

OpenStudy (iimanii):

can somone please just show me the right way to do it this is giving me a headache

pooja195 (pooja195):

Basically they asked you what 34 * 68 is which is incorrect it's not how the problem should have been started out \(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 6/34 = p/68 simply cross multiply \[\huge~\rm~\bf~\frac{ 6 }{ 34 }=\frac{ p }{ 68 }\] 34*p=? 68*6=?

rebeccaxhawaii (rebeccaxhawaii):